What are the 6 trigonometric functions for t = 5 pi/ 6
Answer by stanbon(75887)

What are the 6 trigonometric functions for t = 5 pi/ 6
-------
(5/6)pi is in Q2
The reference angle for (5/6)pi is pi/6
----
sin(5/6 pi) = sin(pi/6) = 1/2
cos(5/6 pi) = -cos(pi/6) = -sqrt(3)/2
tan(5/6 pi) = -tan(pi/6) = -1/sqrt(3)
-----------------------
csc(5/6 pi) = csc(pi/6) = 2
sec(5/6 pi) = -sec(pi/6) = 2/sqrt(3)
cot(5/6 pi) = -cot(pi/6) = -sqrt(3)
=========================================
Cheers,
Stan H.